A) For the reaction

4HCl(g) + O2(g)---->2H2O(g) + 2Cl2(g)

Delta H° = -114.4 kJ and Delta S° = -128.9 J/K

The equilibrium constant for this reaction at 306.0 K is _______?

Assume that Delta H° and DeltaS° are independent of temperature.

B) For the reaction

N2(g) + 3H2(g)---->2NH3(g)

Delta H° = -92.2 kJ and Delta S° = -198.7 J/K

The equilibrium constant for this reaction at 333.0 K is ___?

Assume that Delta H° and Delta S° are independent of temperature.
Answer #1

A)

ΔHo = -114.4 KJ

ΔSo = -128.9 J/K

= -0.1289 KJ/K

T = 306 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -114.4 - 306.0 * -0.1289

ΔGo = -74.9566 KJ

= -74956.6 J

Given:

T = 306 K

use:

ΔGo = -R*T*ln Kc

-74956.6 = - 8.314*306.0* ln(Kc)

ln Kc = 29.4631

Kc = 6.247*10^12

Answer: 6.25*10^12

B)

ΔHo = -92.2 KJ

ΔSo = -198.7 J/K

= -0.1987 KJ/K

T = 333 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -92.2 - 333.0 * -0.1987

ΔGo = -26.0329 KJ

ΔGo = -26032.9 J

Given:

T = 333 K

use:

ΔGo = -R*T*ln Kc

-26032.9 = - 8.314*333.0* ln(Kc)

ln Kc = 9.403

Kc = 1.213*10^4

Answer: 1.21*10^4